OpenStudy (anonymous):

4z+2=z-5 solve and check the solution please

OpenStudy (anonymous):

3z=-7 z=-7/3

OpenStudy (anonymous):

thanks how a6x+2=z-4 sole and check solution

OpenStudy (anonymous):

what are you solving for?

OpenStudy (anonymous):

6x+2=z-4

OpenStudy (anonymous):

are you looking for the solution for z or x though?